OpenStudy (watchmath):

Lets play some game! Create integers starting from 1 using exactly four of 4. You may use some standard math notaton : plus,minus, square root, factorial , multiplication, division. You can only proceed to work with the number \(n\) if you (or someone else) already did the \(n-1\).

OpenStudy (watchmath):

Let me start: \(1=4-4+\frac{4}{4}\)

OpenStudy (anonymous):

\[(\log_{4} 4)^{4^{4}}\]

OpenStudy (watchmath):

Thanks Daniel, please mention what the number is equal to :D

OpenStudy (anonymous):

Lol, I did 1 again: \[2= 4/4 + 4/4\]

OpenStudy (watchmath):

\(3=\frac{4+4+4}{4}\)

OpenStudy (watchmath):

\[4=4\cdot 4^{4-4}\]

OpenStudy (watchmath):

\(5=4+4^{4-4}\) \(6=4+\frac{4+4}{4}\) \(7=(4\cdot \sqrt{4})-\frac{4}{4}\)

OpenStudy (anonymous):

\[8 = \sqrt{4}*\sqrt[4]{4^4}\]

OpenStudy (watchmath):

\(9=4+4+\frac{4}{4}\)

OpenStudy (anonymous):

\[10 = \sqrt{4*4*4}+\sqrt{4}\]

OpenStudy (watchmath):

\[11=\frac{4!}{\sqrt{4}}-\frac{4}{4}\] \[12=4+4+\sqrt{4}+\sqrt{4}\] \[13=\frac{4!}{\sqrt{4}}+\frac{4}{4}\]

OpenStudy (anonymous):

\[14 = 4!*\sqrt{4}/4 + \sqrt{4}\]

OpenStudy (watchmath):

\(15=\frac{44}{4}+4\)

OpenStudy (anonymous):

16=4*4*4/4

OpenStudy (anonymous):

17=(4*4)+(4/4)

OpenStudy (anonymous):

\[18 + \sqrt{4}*\sqrt{4}*4 + \sqrt{4}\]

OpenStudy (anonymous):

\[19 = 4! - 4 - 4/4\]

OpenStudy (anonymous):

\[20 = 4*(4+\log_{4}4) \]

OpenStudy (watchmath):

\(21=4!-4+\frac{4}{4}\)

OpenStudy (anonymous):

4!/4*4-\[\sqrt{4}\]

OpenStudy (watchmath):

Let me rewrite metarock result \(22=\frac{4!}{4}\cdot 4-\sqrt{4}\)

OpenStudy (anonymous):

\[23=4!-\sqrt{4} + 4/4\]

OpenStudy (anonymous):

24=4*4 + 4 + 4

OpenStudy (anonymous):

\[24 = 4*(4*\sqrt{4}-\sqrt{4})\]

OpenStudy (anonymous):

\[25 = 4! + \sqrt{4}-4/4\]

OpenStudy (watchmath):

\[25=4!+\frac{\sqrt{4\cdot 4}}{4}\]

OpenStudy (watchmath):

\[26=4!+\frac{4+4}{4}\]

OpenStudy (watchmath):

\[27=4!+4-\frac{4}{4}\]

OpenStudy (anonymous):

28=4!*4/4 + 4

OpenStudy (watchmath):

\(29=4!+4+\frac{4}{4}\)

OpenStudy (anonymous):

\[30=4!+4+\sqrt{4}\]

OpenStudy (watchmath):

Need to use four of 4 metarock!

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

30\[30=4!+\sqrt{4}+\sqrt{4}+\sqrt{4}\]

OpenStudy (anonymous):

watchmath if i want to chat with u should i open the group chat below?

OpenStudy (watchmath):

sure

OpenStudy (watchmath):

\[31=(\sqrt{\sqrt{\sqrt{4^{4!}}}}/-\sqrt{4} )/\sqrt{4}\]

OpenStudy (anonymous):

too good probably this is the toughest one

OpenStudy (watchmath):

\(32=4\cdot 4 + 4\cdot 4\) \(33=(\sqrt{\sqrt{\sqrt{4^{4!}}}}+\sqrt{4})/\sqrt{4}\)

OpenStudy (watchmath):

\(34=\left (\sqrt{\sqrt{\sqrt{4^{4!}}}} \right )/\sqrt{4}+\sqrt{4}\)

OpenStudy (anonymous):

combinations of this can flit an 320 GB pc memory :P

OpenStudy (watchmath):

You want to continue to play? Try 35

OpenStudy (anonymous):

:P ok i try

OpenStudy (watchmath):

Ah maybe I don't know 35 either :(

OpenStudy (watchmath):

I know it :D

OpenStudy (anonymous):

\[\sum_{4}^{8}n \] mybe try this xd

OpenStudy (watchmath):

But you can't use 8 :)

OpenStudy (anonymous):

wathcmath how old are you?

OpenStudy (watchmath):

very old so don't ask :D :D

OpenStudy (anonymous):

4+4 =8 i tihnk i can use xd

OpenStudy (anonymous):

just say xd

OpenStudy (watchmath):

Ok, but then you run out of your 4

OpenStudy (anonymous):

no :( there are 4 fours ^^

OpenStudy (anonymous):

and I am 14 years old ^^or we can say 13,9911

OpenStudy (anonymous):

13,8911*

OpenStudy (watchmath):

How's life in Turkey?

OpenStudy (anonymous):

really nice :) i am happy in there ^^

OpenStudy (anonymous):

ad nd i lreally liked that site :) i learned limit and multiply today ^^ with the help of this site ^^

OpenStudy (watchmath):

you can visit my site too and tell me what you think http://ask.watchmath.com

OpenStudy (anonymous):

okay but how old are you ?

OpenStudy (anonymous):

:D :D

OpenStudy (nowhereman):

Ok, maybe you see this as cheating, but I wrote a short python script (46lines) that does find these things with at most 4 4's. I limited it to intermediate results of 100 so I got none for 33. But for 31 there's a nicer one: \[31 = 4! + \frac{4!+4}{4}\]

OpenStudy (watchmath):

wow! that's good nowhereman I wish I can do programming.

OpenStudy (anonymous):

WIOW!

OpenStudy (watchmath):

I want to learn programming, is python a good way to start?

OpenStudy (nowhereman):

It's easy to use and has many advances features, but I think for the beginner it is not strict enough. For example you have very loose typing of variables. On the other hand the syntax is really nice. I'll just append the script, if you want to have a look :-)

OpenStudy (watchmath):

Are you doing your diplom right now nowhereman? typically how long does it takes?

OpenStudy (nowhereman):

Here in germany, it normally takes 9 semester. But I'm one of the last ones to do a diplom. Now they all do bachelors/masters which take 6+4 semester mostly.

OpenStudy (watchmath):

Then how much longer to get a doctorate? It seems you are interested in functional analysis :)

OpenStudy (nowhereman):

Anything up from 2 years I guess; but I'm not completely sure if I want to go that path yet :-)

OpenStudy (watchmath):

How many more years for you to finish your diplom?